OpenStudy (anonymous):

A furniture rental company charges a base rate plus a rate per day for an office furniture set. There are two available plans. The super saver plan charges $100 + $5 per day. The best deal plan charges $150+$3 per day. How many days make the super saver plan more expensive than the best deal plan? SHOW ALGEBRAIC INEQUALITY

OpenStudy (anonymous):

Let x be the number of days you rent for. You want to know when\[Plan \ A > Plan\ B\]

OpenStudy (anonymous):

If Plan A = 100 + 5x and Plan B = 150 + 3x. You plug in those expressions and solve the inequality.

OpenStudy (anonymous):

A=100+5x>150+3x 2x<50 ? x<25 x>25 Plan A is more expensive after the 25th day the 26th day ?

OpenStudy (anonymous):

You only change the direction of the inequality if you divide or multiply by 0.

OpenStudy (anonymous):

A=100+5x>150+3x 2x>50 x>25 x>25 Plan A is more expensive after the 25th day the 26th day ?

OpenStudy (anonymous):

O k thanks I keep forgetting I need to know if I am writing it out correctly

OpenStudy (anonymous):

If x > 25 then the number of days you rented must be greater than 25.

OpenStudy (anonymous):

oops should that be 20?

OpenStudy (anonymous):

for plan a to be more expensive.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

You wrote it correctly. Plan A is more expensive after the 25th day.

OpenStudy (anonymous):

and the 25 is right ? its not 20

OpenStudy (anonymous):

No. if 2x > 50 than x > 50/2 and 50/2 is 25